# Know Your Needs Better: Towards Structured Understanding of Marketer Demands with Analogical Reasoning Augmented LLMs

### v1

@@ 3.2.1. Knowledge Distillation
We collect nearly 10,000 demands from our daily marketing environment. Following the analogical reasoning method mentioned above, we construct input prompts for each demand in the form of Equation (2), where analogical examples come from the previously constructed reasoning library. Subsequently, based on these prompts that incorporate analogical knowledge, we call the gpt-3.5-turbo API to generate SELL answers, which will then be further verified and corrected through manual crowdsourcing. We denote this part of data comes from demand to answer paradigm.
@@
$\bullet$ Answer generation: As shown in Table 5, we first construct a series of logical templates of SELL, where c is the basic conditional expression in SELL. When constructing an answer, we first randomly select an answer template from the logical template table and then construct the content of every c in the selected template to obtain a synthesized answer. The content of c is obtained from the marketing database through random sampling. Specifically, we first randomly sample a tag from the database as a key in c. Then we sample its operator based on the type of keys. Finally, we randomly select a value from all possible values of that key, thus constructing a c in the form of (key#operator#value). To ensure that the synthesized answers are logically reasonable, we will limit the number of conditional expressions in the answer template, and abandon overly complex intersection and union operations.
$\bullet$ Demand generation: Considering the strong writing ability of ChatGPT, we provide some ($d_{e},s_{e}$) examples in $\mathcal{R}$ as demonstrations to guide gpt-3.5-turbo to write demands given above synthesized answers:
(3) $concat((d_{e},s_{e}),s_{i})\xrightarrow{gpt-3.5-turbo}d_{i}$
where $s_{i}$ is the above synthesized SELL answers. We will check the language fluency and logical rationality of the generated demand, and the specific prompt content is shown in Appendix C.
@@ 3.2.2. Reasoning-Augmented Multi-Task Fine-tune
The training objective of the normal distillation method is to minimize the loss L:
@@
It (Hsieh et al., 2023; Kim et al., 2023) has been pointed out that fine-tuning LLMs using data combined with reasoning steps can enhance the reasoning capabilities of the LLMs. Inspired by (Hsieh et al., 2023), we first prompt gpt-3.5-turbo to generate intermediate reasoning steps of $\mathcal{D}_{train}$ based on reasoning library $\mathcal{R}$:
(5) $concat(({d}_{e},r_{e},s_{e}),(d_{i},s_{i}))\xrightarrow{gpt-3.5-turbo}r_{i}$
@@
The final loss function of multi-task fine-tuning is then defined as:
@@ 4.1.1. Benchmarks
We manually construct a NL2SELL testing benchmark for predicting, which contains 170 pieces of data based on expert knowledge in the format $\mathcal{D}_{test}=\{(d_{p},s_{p})\}_{p=1}^{170}$. The demands in $\mathcal{D}_{test}$ do not exist in the training set $\mathcal{D}_{train}$ or reasoning library $\mathcal{R}$, eliminating the issue of label leakage. Table 1 shows the statistics of these 3 parts of data [^3] The dataset does not contain any Personal Identifiable Information (PII). The dataset is desensitized and encrypted. Adequate data protection was carried out during the experiment to prevent the risk of data copy leakage..
@@ 4.1.2. Baseline
We first use ChatGPT to test the effectiveness of the analogical reasoning based prompting method. For comparison, we propose four additional variants of prompts as baselines:
$\bullet$ Zero-shot: Only given the instruction $\mathcal{I}$ and a test demand $d_{p}$.
$\bullet$ Fixed few-shot: Given the instruction $\mathcal{I}$, k fixed examples $\{(d_{e},s_{e})\}_{e=1}^{k}\in\mathcal{R}$ for all $d_{p}\in\mathcal{D}_{test}$ and a $d_{p}$.
$\bullet$ Fixed few-shot + RS (Reasoning Steps): Given the instruction $\mathcal{I}$, k fixed examples $\{(d_{e},r_{e},s_{e})\}_{e=1}^{k}\in\mathcal{R}$ for all $d_{p}\in\mathcal{D}_{test}$ and a $d_{p}$.
$\bullet$ Random few-shot + RS: Given the instruction $\mathcal{I}$, k randomly sampled examples $\{(d_{e},r_{e},s_{e})\}_{e=1}^{k}\in\mathcal{R}$ for every $d_{p}\in\mathcal{D}_{test}$ and a $d_{p}$. Different $d_{p}$ has different randomly sampled examples.
@@ 4.1.3. Evaluation metrics
$\bullet$ Human evaluation (Human Eval.): We conduct more fine-grained evaluations through human evaluation. Annotators who have received good training will score the accuracy of the key, value, and logic between 0 and 10, and finally provide an overall score. The accuracy of the key and value refers to whether the key and value in predicted answers are identical to or synonymous with labels. The accuracy of logic refers to whether the logic aligns with labels. Human evaluation is important because sometimes the model output may not be completely consistent with labels, but it may also be correct. For example, “(Age#Between#18,35)” could also be expressed as “(Age#Greater Than#18) AND (Age#Less Than#35)”, while the above evaluation metrics would treat them as different answers.
$\bullet$ GPT4 evaluation (GPT4 Eval.): We use the most advanced LLM “GPT4” as the judges. To be specific, we prompt the GPT4 to score the predicted answers from 0 to 10 and provide different scoring examples as a reference for GPT4 (as shown in Appendix D).
@@ 4.1.4. Implementation details
Considering the limitations of input length, we set the number of analogical examples k as 2. In the fine-tuning stage, ChatGLM2-6B-32K and Baichuan2-13B-Chat are tuned with LoRA (Hu et al., 2021) on 1 and 4 A100 (80G). We set the optimizer, maximum context size, batch size, and learning rate to Adam, 4096, 8, and 5e-5 respectively. For testing, a single 80G A100 is used. We use the code framework from LLaMA-Factory [^4] https://github.com/hiyouga/LLaMA-Factory (hiyouga, 2023).
Table 2. NL2SELL results on ChatGPT. The best results are in bold, while the second are underlined.
@@
ARAP 57.0 0.824 0.824 0.807 0.653 0.715 0.678 0.774 6.84
4.2. Analogical Prompting result on ChatGPT
@@ 4.2.1. Main Results
Table 2 displays the results of different prompting methods on ChatGPT, where ARAP represents our analogical reasoning augmented prompting. The experimental results demonstrate that: (1) ARAP significantly outperforms other prompting methods in various evaluation metrics, especially in terms of structural accuracy which improves by over 7%, proving that analogical examples provide a good logical structure for reasoning. (2) Even randomly selected few-shot reasoning examples are better than fixed ones. This indicates that fixed reasoning examples for all inputs are often suboptimal, demonstrating the necessity of establishing a reasoning library to provide diverse reasoning samples. (3) Reasoning steps are essential in prompting. When comparing results between “Fixed few-shot” and “Fixed few-shot + RS”, we found that reasoning steps have a significant active impact on reasoning accuracy, which is similar to the trend in other previous research works (Wei et al., 2022).
@@ 4.2.2. Case Study
Table 6 shows some cases we obtain from ChatGPT with different versions of prompts. Although we provide the same tag list for a specific demand across different prompt versions, there are significant variations in the selection of key-value pairs. For the first demand “Company white-collar workers who enjoy drinking Starbucks”, all baseline versions ignore the tag of “Career”, indicating that the model cannot smoothly map the description “white-collar workers” to “Career”, although this is common sense. Meanwhile, in both fixed few-shot versions, the model treats “company white-collar workers” as the value of the “Resident City”, which is due to the answers of fixed examples in prompts containing similar expressions such as (Resident city#Belongs To#Hangzhou), resulting in the model being inappropriately imitated. While in ARAP, we retrieve similar demands like “Young people who enjoy swimming or female white-collar workers who enjoy reading in Shanghai” and its reasoning steps from the reasoning library, thus helping the LLM analogically transform the “white-collar workers” into conditional expression “(Career#Belongs To#White-collar)” as they share the similar description “white-collar workers”. In another case “Young men who enjoy collecting toy models”, there are many grammar errors and inappropriate expressions in baseline answers, like non-existent operators “Contain” and inappropriate value “Young” for key “Age”. More details are described in the Appendix B.
@@ 4.3.1. Main Results
Table 3 shows the testing results on fine-tuned LLMs with the knowledge distilled from ChatGPT, where ARAFT represents fine-tuning (FT) in the reasoning-augmented multi-task paradigm mentioned in section 3.2.2. All testing prompts are the same as ARAP prompts in ChatGPT. As is shown: (1) Both two models show comparable capability to ChatGPT, especially on Baichuan2-13B-Chat, where the score of S-BLEU is improved by over 5% compared to ChatGPT after fine-tuning. This demonstrates the superiority of our distillation and training methods. (2) Compared with the results of the single-task training strategy (-MT), the multi-task training strategy brings improvement to the reasoning performance of the LLMs. Although during testing, we limit the model to only output the final answers without reasoning steps, adding training tasks that predict reasoning steps can enhance the models’ reasoning abilities and robustness. (3) Compared with the results without reasoning steps in the input corpus (-RS), we can conclude that providing explicit reasoning steps in training input is beneficial. If there are no reasoning steps in the input corpus but a multi-task training strategy is used, it will increase the difficulty of the training task and lead to poor performance, which is particularly evident in Baichuan2-13B-Chat. (4) The result of fine-tuning using only the normal demand and answer pairs is not satisfactory, which indicates that it is challenging for LLMs to learn the data patterns solely based on the Q&A pairs in situations where tasks are difficult. Therefore, it is necessary to design appropriate demonstrations and training tasks to fine-tune the LLMs.
@@ 5. Application
Application Case. Figure 3 shows an application case of ARALLM in online user targeting. When the marketers input their demand and click the “Search” button, our system will invoke the retriever service to retrieve the demands from the reasoning library most related to the current demand for generating the prompt based on analogical prompting and request the deployed ARAFT model to generate the answer expressed in SELL. The SELL expression will be parsed by our parser and results will be visualized to the marketers on the panel. The marketers can verify and edit the card on the panel (Figure 3) after clicking the “Edit” button, such as modifying the operator “AND” to “OR”. Finally, they can click the “Export” button to get the target users. Our system has been already running for months. We have set two metrics to evaluate the new approach online, including operation time, and number of likes from marketers. By collecting the marketers’ operation log and questionnaire surveys, we find the new system’s operation time is 4-10 times shorter than the former on average, and the number of likes is 1.3 times larger than the former, revealing the advantage of this new way of user targeting.
@@ 7. Conclusion and future work
In this paper, we provide a novel user targeting approach by leveraging LLMs to gain a structured understanding of marketers’ demands. We first define a new language SELL to express the needs of marketers more clearly. Subsequently, we propose an analogical reasoning augmented framework, ARALLM, which consists of analogical reasoning based prompting and reasoning-augmented multi-task model distillation. The experimental results on ChatGPT show that our analogical reasoning based prompting significantly outperforms other baseline prompting methods in the NL2SELL task. In addition, we distill a large-scale dataset using ChatGPT and train the student LLMs using a multi-task training approach, which is successfully used for online deployment.
@@
Appendix A Logic template
Table 5 shows the logic template we use for answer generation in section 3.2.1. The c in the template represents the basic conditional expression in SELL.

### v3

@@ 2.3.1. Knowledge Distillation
We collect nearly 10,000 demands from our daily marketing environment. Following the analogical reasoning method mentioned above, we construct input prompts for each demand in the form of Equation (2), where analogical examples come from the previously constructed reasoning library. Subsequently, we call the GPT-3.5 API to generate SELL answers based on the above prompts, which will then be further verified and corrected through manual crowdsourcing. We denote this part of data comes from demand to answer paradigm.
@@
$\bullet$ Answer generation: We first construct a series of logical templates of SELL. The content of logical templates is obtained from the marketing database through random sampling. Specifically, we first randomly sample a tag from the database as a key. Then we sample its operator based on the type of keys. Finally, we randomly select a value from all possible values of that key, thus constructing a unit conditional expression in the form of (key#operator#value). Different sampled conditional expressions can be combined by union and intersection to form different types of targeted users. To ensure that the synthesized answers are logically reasonable, we will limit the number of conditional expressions in the logical template, and abandon overly complex intersection and union operations.
$\bullet$ Demand generation: Considering the strong ability of GPT-3.5, we provide some ($d_{e},s_{e}$) examples in $\mathcal{R}$ as demonstrations to guide GPT-3.5 to write demands given above synthesized answers:
(3) $concat((d_{e},s_{e}),s_{i})\xrightarrow{GPT-3.5}d_{i}$
where $s_{i}$ is the synthesized SELL answers. The language fluency and logical rationality of the generated demand will be checked.
@@ 2.3.2. Reasoning-Augmented Multi-Task Fine-tune
The training objective of the normal distillation is to minimize the loss L:
@@
It (Hsieh et al., 2023; Kim et al., 2023) has been pointed out that fine-tuning LLMs using data combined with reasoning steps can enhance the reasoning capabilities of the LLMs. Inspired by (Hsieh et al., 2023), we first prompt GPT-3.5 to generate intermediate reasoning steps of $\mathcal{D}_{train}$ based on $\mathcal{R}$:
(5) $concat(({d}_{e},r_{e},s_{e}),(d_{i},s_{i}))\xrightarrow{GPT-3.5}r_{i}$
@@
The final loss function of multi-task fine-tuning is defined as:
@@ 3.1.1. Benchmarks
We manually construct an NL2SELL testing benchmark for predicting, which contains 170 pieces of data based on expert knowledge in the format $\mathcal{D}_{test}=\{(d_{p},s_{p})\}_{p=1}^{170}$. The demands in $\mathcal{D}_{test}$ do not exist in the training set $\mathcal{D}_{train}$ or reasoning library $\mathcal{R}$, eliminating the issue of label leakage. Table 1 shows the statistics of these 3 parts of data [^1] The dataset does not contain any Personal Identifiable Information (PII). The dataset is desensitized and encrypted. Adequate data protection was carried out during the experiment to prevent the risk of data copy leakage.. It should be noted that all training data generated by GPT-3.5 are strictly verified by well-trained volunteers, and we provide some examples in the supplementary materials. We will open source all these datasets on request.
@@ 3.1.2. Baseline
We first use GPT-3.5 to test the effectiveness of the analogical reasoning based prompting method. For comparison, we propose four additional variants of prompts as baselines:
$\bullet$ Zero-shot: Only given the instruction $\mathcal{I}$, tag list $T_{p}$ and a test demand $d_{p}$.
$\bullet$ Fixed few-shot: Given the instruction $\mathcal{I}$, k fixed examples $\{(d_{e},s_{e})\}_{e=1}^{k}\in\mathcal{R}$ for all $d_{p}\in\mathcal{D}_{test}$, tag list $T_{p}$ and a $d_{p}$.
$\bullet$ Fixed few-shot + RS (Reasoning Steps): Given the instruction $\mathcal{I}$, k fixed examples $\{(d_{e},r_{e},s_{e})\}_{e=1}^{k}\in\mathcal{R}$ for all $d_{p}\in\mathcal{D}_{test}$, tag list $T_{p}$ and a $d_{p}$.
$\bullet$ Random few-shot + RS: Given the instruction $\mathcal{I}$, k randomly sampled examples $\{(d_{e},r_{e},s_{e})\}_{e=1}^{k}\in\mathcal{R}$ for every $d_{p}\in\mathcal{D}_{test}$ ,tag list $T_{p}$ and a $d_{p}$. Different $d_{p}$ has different randomly sampled examples.
@@ 3.1.3. Evaluation metrics
$\bullet$ Human evaluation (Human Eval.): We conduct more fine-grained evaluations through human evaluation. Annotators who receive good training will score the accuracy of the key, value, and logic between 0 and 10, and finally provide an overall score.
$\bullet$ GPT4 evaluation (GPT4 Eval.): We use the most advanced LLM gpt-4-turbo-preview (GPT-4) as the judges. To be specific, we prompt the GPT-4 to score the predicted answers from 0 to 10 and provide different scoring examples as a reference for GPT-4.
@@ 3.1.4. Implementation details
We set the number of analogical examples k as 3. The size of the reasoning library $\mathcal{N}_{RL}$ is 100. In the fine-tuning stage, ChatGLM2-6B-32K and Baichuan2-13B-Chat are tuned with LoRA (Hu et al., 2021) on 1 and 4 A100 (80G). We set the optimizer, maximum context size, batch size, and learning rate to Adam, 4096, 8, and 5e-5 respectively. For testing, a single 80G A100 is used. We use the code framework from LLaMA-Factory [^2] https://github.com/hiyouga/LLaMA-Factory (hiyouga, 2023).
Table 2. NL2SELL results on GPT-3.5. The best results are in bold, while the second are underlined.
@@
ARAP 58.5 0.779 0.809 0.794 0.653 0.715 0.678 0.774 6.84
3.2. Analogical Prompting result on GPT-3.5
@@ 3.2.1. Main Results
Table 2 displays the results of different prompting methods on GPT-3.5, where ARAP represents our Analogical Reasoning Augmented Prompting. The experimental results demonstrate that: (1) ARAP significantly outperforms other prompting methods in various evaluation metrics, especially in terms of structural accuracy which improves by over 7%, proving that analogical examples provide a good logical structure for reasoning. (2) Even randomly selected few-shot reasoning examples are better than fixed ones. This indicates that fixed reasoning examples for all inputs are often suboptimal, demonstrating the necessity of establishing a reasoning library to provide diverse reasoning samples. (3) Reasoning steps are essential in prompting. When comparing results between “Fixed few-shot” and “Fixed few-shot + RS”, we found that reasoning steps have a significant active impact on reasoning accuracy, which is similar to the trend in other previous research works (Wei et al., 2022).
Figure 3. The impact of the size of parameter settings.
Further, we explore the impact of parameter settings, such as the size of reasoning library $\mathcal{N}_{RL}$ and the number of analogical examples $k$. As shown in Figure 3 (a), as the size of the reasoning library increases, the overall accuracy gradually increases, but the rate slows down, indicating that when the inference library reaches a certain scale, it can already provide references for most marketer’s demands. In Figure 3 (b), it can be observed that the overall accuracy first increases and then decreases when the number of examples in prompts increases. This reflects that when engaging in context-based learning, providing more examples to LLMs is not always better. When an excessive number of examples are input, irrelevant examples may introduce noise, and an overly lengthy context can also impose a greater burden on the model’s understanding.
@@ 3.2.2. Case Study
Table 3. Case study on GPT-3.5 with different versions of prompt. Bad responses are marked with a wavy underline.
Demand:Company white-collar workers who enjoy drinking Starbucks
Zero-shot (Preference#Belongs To#Starbucks)
Fixed few-shot (Preference#Belongs To#Starbucks) AND (Resident City#Belongs To#Company white-collar)
Fixed few-shot + RS (Preference#Belongs To#Enjoy drinking Starbucks) AND (Resident City#Belongs To#Company white-collar)
Random few-shot + RS (Preference#Belongs To#Starbucks)
Table 3 is a case we obtained from GPT-3.5 with different versions of prompts. Although we provide the same tag list for a specific demand across different prompt versions, there are significant variations in the selection of key-value pairs. For the first demand “Company white-collar workers who enjoy drinking Starbucks”, all baseline versions ignore the tag of “Career”, indicating that the model cannot smoothly map the description “white-collar workers” to “Career”, although this is common sense. Meanwhile, in both fixed few-shot versions, the model treats “company white-collar workers” as the value of the “Resident City”, which is due to the answers of fixed examples in prompts containing similar expressions such as (Resident city#Belongs To#Hangzhou), resulting in the model being inappropriately imitated. While in ARAP, we retrieve similar demands like “Young people who enjoy swimming or female white-collar workers who enjoy reading in Shanghai” and its reasoning steps from the reasoning library, thus helping the LLM analogically transform the “white-collar workers” into conditional expression “(Career#Belongs To#White-collar)” as they share the similar description “white-collar workers”.
@@ 3.3.1. Main Results
Table 4 shows the testing results on fine-tuned LLMs with the knowledge distilled from GPT-3.5, where ARAFT represents fine-tuning (FT) in the reasoning-augmented multi-task paradigm mentioned in section 2.3.2. All testing prompts are the same as ARAP prompts in GPT-3.5. As is shown: (1) Both two models show comparable capability to GPT-3.5, especially on Baichuan2-13B-Chat, where the score of S-BLEU is improved by over 5% compared to GPT-3.5 after fine-tuning. This demonstrates the superiority of our distillation and training methods. (2) Compared with the results of the single-task training strategy (-MT), the multi-task training strategy brings improvement to the reasoning performance of the LLMs. Although during testing, we limit the model to only output the final answers without reasoning steps, adding training tasks that predict reasoning steps can enhance the models’ reasoning abilities and robustness. (3) Compared with the results without reasoning steps in the input corpus (-RS), we can conclude that providing explicit reasoning steps in training input is beneficial. If there are no reasoning steps in the input corpus but a multi-task training strategy is used, it will increase the difficulty of the training task and lead to poor performance, which is particularly evident in Baichuan2-13B-Chat. (4) The result of fine-tuning using only the normal demand and answer pairs is not satisfactory, which indicates that it is challenging for LLMs to learn the data patterns solely based on the Q&A pairs in situations where tasks are difficult. Therefore, it is necessary to design appropriate demonstrations and training tasks to fine-tune the LLMs.
@@ 4. Application
Application Case. Figure 4 shows an application case of ARALLM in online user targeting. A marketer wants to do marketing for a stage drama performance, so he can input a raw demand and click the “Search” button. Our system will invoke the retriever service to retrieve similar top-k demonstrations and top-n tags from the reasoning library embeddings (RL Embs) and tag embeddings (Tag Embs), respectively. After filling the prompt based on retrieved demonstrations and tags, our system requests the deployed ARAFT model to generate the answer expressed in SELL. The SELL expression will be parsed by our parser and results will be visualized to the marketers on the panel as shown in Figure 4. The marketers can verify and edit the card on the panel after clicking the edit button. Finally, they can click the “Export” button to get the target users.
Practical Effects. Our system has been already running for months. We have set two metrics to evaluate the new approach online, including operation time, and number of likes from marketers. By collecting the operation log, the entire system takes no more than 10 seconds to complete a request and the average marketers’ operation time of the new system is about 3 minutes, which is 4-10 times shorter than the former on average. Beyond that, we receive likes 1.3 times larger than the former, revealing the operation friendliness of this new way of user targeting.
Table 6. Online A/B testing results.
Campaign CTR Exposure
Digital car +55.21% +0.33%
Housing provident fund +30.56% -0.21%
Ophthalmic health +69.17% +0.01%
Traditional Chinese Medicine +112.7% +0.67%
Recharge +73.89% +0.20%
Online A/B testing results. We conduct online A/B experiments to compare our methods with the traditional rule-based methods and use the direct metrics CTR (Click-Through-Rate) and Exposure (the number of users who have been exposed by the campaign) to evaluate them. Results are shown in Table 6. From an end-to-end perspective, our solution has achieved better marketing results than rule-based methods in multiple marketing campaigns.
@@ 6. Conclusion and future work
In this paper, we provide a novel user targeting approach by leveraging LLMs to gain a structured understanding of marketers’ demands. We first define a new language SELL to express the needs of marketers more clearly. Subsequently, we propose an analogical reasoning augmented framework, ARALLM, which consists of analogical reasoning based prompting and reasoning-augmented multi-task model distillation. The experimental results on GPT-3.5 show that our analogical reasoning based prompting significantly outperforms other baseline prompting methods in the NL2SELL task. In addition, we distill a large-scale dataset using GPT-3.5 and train the student LLMs using a multi-task training approach, which is successfully used for online deployment.
@@
Acknowledgements.
This work is funded by NSFC62306276, NSFCU23B2055 and NSFCU19B2027. This work is supported by the Fundamental Research Funds for the Central Universities (226-2023-00138). This work was supported by Ant Group.
Appendix A Design of SELL
We provide a detailed explanation of the design and the usage of the SELL. The SELL is mainly composed of four elements:
Keys are a series of tags that describe the features of the user, such as Gender, Marital Status, Resident City, and so on. In practical marketing scenarios, the database contains tens of thousands of tags that describe user profiles. Selecting the appropriate keys from this vast array of tags is a challenging task.
Values are the fillings of the corresponding keys. They can be generally divided into three types: numerical type, string type, and boolean type. For instance, the value corresponding to the key Monthly Income is of numerical type, because they are continuous and non-enumerable. The value of Gender is of string type, with limited options: Male or Female. Boolean type values have only two states, i.e., True or False, such as Marital Status.
Operators represent the relationship between keys and values. For numerical values, seven types of operators can be used to connect them with corresponding keys: Equal To, Greater Than, Less Than, Not Equal To, Not Greater Than, Not Less Than, and Between. For values of string type and boolean type, there are two types of operators: Belongs To and Not Belongs To. Keys, operators, and values can form a basic conditional expression in SELL, formatted as “(key#operator#value)”, such as “(Marital Status#Belongs To#True)”. A conditional expression represents a cluster of target users, for example, in the above example, the target users are married.
Intersection and Union symbols are used to take the intersection or union of multiple conditional expressions. In practical marketing scenarios, the features of the target users are often complex and difficult to describe using a single conditional expression. Therefore, it is necessary to combine the target users through intersection or union operations. In SELL, we use the symbols AND and OR to denote the intersection and union operations, respectively. For example, “((Resident City#Belongs To#City A) OR (Resident City#Belongs To#City B)) AND (Pet Owning#Belongs To#True)” describes the pet owners who live in City A or City B.
Some demands and their corresponding SELL expressions are shown in Table 7. More data can be found at https://github.com/alipay/Analogic-Reasoning-Augmented-Large-Language-Model.
Appendix B Human-Written Reasoning Steps
As mentioned in the main paper, we write reasoning steps for 10% of data in the reasoning library. Table 8 shows an example.
Appendix C Reasoning Steps Generation Prompt
As mentioned in the main paper, we use GPT-3.5 to help us complete the reasoning steps of remained data. Figure 8 shows the prompt we use in this step.
Figure 5. Reasoning steps generation prompt.
Figure 6. Demand generation prompt.
Table 7. Some examples of marketers’ demands and corresponding SELL.
Middle-aged women (User Age Group#Between#35,55) AND (Gender#Belongs To#Female)
Young people in City A who enjoy swimming or female white-collar workers in City B who enjoy reading ((Resident City#Belongs To#City A) AND (Preference#Belongs To#Swimming) AND (User Age Group#Between#18,35)) OR ((Resident City#Belongs To#City B) AND (Gender#Belong To#Female) AND (Career#Belongs To#White-collar) AND ((Preference#Belongs To#Reading) OR (Days of Listening To Audiobooks#Greater Than#1)))
Users who have a certain understanding or interest in financial products, insurance, etc. (Preference#Belongs To#Wealth Management Products) OR (Preference#Belongs To#Insurance) OR (Eligible group for Wealth Infinity Card#Belongs To#True) OR (Has actively invested in major financial products#Belongs To#True)
Table 8. Examples of human-written reasoning steps.
Demand Mothers who pay attention to baby education.
Tag List User Has Child/Days of Listening to Audiobooks/Homepage Visits/Red Packet Cover Collection/Current Insured Products/Balance of Points/Alipay User in Gourmet Scene/User Age Group/Number of Days Since User’s First Account Opening/User Child Age/User Gender/User Marital Status/Preferences/……
Reasoning Steps (1)Extract keywords: Mothers, baby education (2) Select tags: User Has Child, User Child Age, User Gender, User Marital Status, Preferences (3) Form conditional expressions: (User Has Child#Belongs To#True) (User Gender#Belongs To#Female) (User Child Age#Between#0,4) (Preference#Belongs To#Baby Education) (4) Combine: (Preference# Belongs To# Baby Education) AND (User Gender#Belongs To# Female) AND ((User Has Child#Belongs To#True) OR (User Child Age#Between#0,4))
Answers (Preference#Belongs To#Baby Education) AND (User Gender#Belongs to#Female) AND ((User Has Child#Belongs To#True) OR (User Child Age#Between#0,4))
Figure 7. GPT4 Evaluation prompt.
Figure 8. Reasoning steps generation prompt.
Appendix D Demand generation prompt
Appendix E Logic Template
As mentioned in the main paper, we first construct a series of logical templates of SELL to artificially synthesize answers. Table 9 shows all the logical templates we used. The c in the template represents the basic conditional expression in SELL. To ensure that the synthesized answers are logically reasonable, we limit the number of conditional expressions in the logical template, and abandon overly complex intersection and union operations.